# Thermal quasi-particle theory

### v1

@@ V Thermodynamic functions
$10^{5}$ $-101.02137$ $-103.48646$ $-101.30273$ $-102.10659$
$10^{6}$ $-150.56294$ $-151.43748$ $-150.56368$ $-151.24440$
$10^{7}$ $-729.93806$ $-730.10421$ $-729.93862$ $-730.09519$
$10^{8}$ $-6846.98049$ $-6847.00261$ $-6846.98055$ $-6847.00247$
Figure 3: The deviation from thermal FCI theory in grand potential $\Omega$ of the hydrogen fluoride molecule (0.9168 Å) in the STO-3G basis set as a function of temperature. The “HF” stands for thermal HF theory [thermal QP(1) theory], “MBPT(2)” denotes the second-order finite-temperature many-body perturbation theory with the zero-temperature HF reference, “QP(2)” designates the second-order thermal quasi-particle theory with the zero-temperature HF reference.
@@
$10^{4}$ $-98.57076$ $-98.58809$ $-98.58810$ $-98.59658$
$10^{5}$ $-97.94385$ $-97.86604$ $-97.94314$ $-98.04938$
$10^{6}$ $-96.79410$ $-96.99284$ $-96.78681$ $-96.94534$
$10^{7}$ $-92.02773$ $-92.05724$ $-92.02752$ $-92.05557$
$10^{8}$ $-88.48266$ $-88.48744$ $-88.48268$ $-88.48740$
@@
Tables 14 list the thermodynamic functions ($\Omega$, $U$, $\mu$, and $S$) of thermal HF [thermal QP(1)],Gu and Hirata (2024) thermal QP(2), fintie-temperature MBPT(2),Hirata and Jha (2020); Hirata (2021a) and thermal FCI theoriesKou and Hirata (2014) for an ideal gas of the identical hydrogen fluoride molecules in a wide range of temperature ($T$). Figures 36 plot the deviations from thermal FCI benchmarks, the latter being exact within a basis set. The thermal QP(2) and finite-temperature MBPT(2) calculations were based on the zero-temperature HF reference. See Ref. Gu and Hirata, 2024 for the data for the same system from Fermi–Dirac [thermal QP(0)] theory and other thermal mean-field theories.
In all cases, thermal QP(2) results are close to finite-temperature MBPT(2) ones at low $T$ and are more accurate than thermal HF theory. This is expected because both thermal QP(2) theory and finite-temperature MBPT(2) reduce to zero-temperature MBPT(2), accounting for electron correlation, while thermal HF theory and its zero-temperature counterpart do not. At higher $T$, thermal QP(2) theory quickly approaches thermal HF theory, where their quasi-independent-particle picture becomes relatively less erroneous. However, at $T=10^{6}$ K and $10^{7}$ K, finite-temperature MBPT(2) is systematically more accurate as it accounts for the effect of electron correlation in $S$ and $\mu$ and it uses more complete second-order corrections (including anomalous-diagram terms) in $\Omega$ and $U$. At the intermediate temperature of $T=10^{5}$ K, thermal QP(2) theory (or even thermal HF theory) outperforms finite-temperature MBPT(2). This may be ascribed to the fact that thermal QP(2) theory includes the effect of electron correlation in the orbital energies, and thus higher-order correlation corrections in a manner similar in spirit to self-consistent Green’s function methods,Luttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Van Neck et al. (1991); Dickhoff (1999); Dickhoff and Barbieri (2004); Dahlen and van Leeuwen (2005); Barbieri (2006); Barbieri and Dickhoff (2009); Phillips and Zgid (2014); Neuhauser et al. (2017); Tarantino et al. (2017); Kadanoff and Baym (2018); Coveney and Tew (2023) which are said to work for strong correlation. However, this similarity must not be overstated because self-consistent Green’s function methods replace reference orbitals and orbital energies by electron-correlated counterparts, while thermal QP(2) theory uses correlated orbital energies only in the Fermi–Dirac distribution functions and thus reduces identically to MBPT(2) at $T=0$.
VI Thermal quasi-particle energies
@@
$\epsilon_{p}\equiv\frac{\partial U}{\partial f_{p}^{-}},$ (48)
encompassing Fermi–Dirac, thermal HF, and thermal QP(2) theories.
The $\epsilon_{p}$ has the literal physical meaning of the increase in the internal energy upon infusion of an infinitesimal fraction of an electron in the $p$th spinorbital. While it is not irrational to speak about a fraction of an electron here because the process in question is the thermal average of the same processes involving an infinite number of particles, it is unclear how this quantity can be used in relation to observables. The $\epsilon_{p}$ at $T=0$ is rigorously connected with an ionization or electron-attachment energy via Koopmans’ theoremSzabo and Ostlund (1982) or MBGF.Hirata et al. (2017) However, such an interpretation will face difficulty when $T\gg 0$ because ionization or electron attachment at finite temperature will involve different orbitals at different probabilities, even in a strict independent-particle picture.
Just for convenience of the following analysis, let us define the internal energy $U(N)$ as a function of the average number of electrons $N$ (which may now differ from $\bar{N}$) by the respective formula of $U$ with only the chemical potential adjusted. For instance, the internal energy of thermal FCI theory is given by
$U^{\text{FCI}}(N)$ $=$ $\frac{\sum_{I}E_{I}^{\text{FCI}}e^{-\beta(E_{I}^{\text{FCI}}-\mu N_{I})}}{\sum_{I}e^{-\beta(E_{I}^{\text{FCI}}-\mu N_{I})}},$ (49)
in which only the $\mu$ is adjusted for a given $N$ to satisfy,
$N$ $=$ $\frac{\sum_{I}N_{I}e^{-\beta(E_{I}^{\text{FCI}}-\mu N_{I})}}{\sum_{I}e^{-\beta(E_{I}^{\text{FCI}}-\mu N_{I})}}.$ (50)
@@
This is a valid and in fact, exact (within a basis set) treatment of thermodynamics if the particles are not electrically charged. For electrons, however, it describes a physically unrealistic, massively charged plasma, whose energy is not even extensive.Fisher and Ruelle (1966); Dyson and Lenard (1967); Hirata and Ohnishi (2012) Nevertheless, $U(N)$ is computationally well defined even for electrons thanks to the ideal-gas assumption (i.e., no interactions between molecules) and will be useful for the following analysis.
For Fermi–Dirac, thermal HF, and thermal QP(2) theories, the internal energy $U(N)$ as a function of $N$ is defined by Eqs. (1), (13), and (LABEL:QP:U), respectively, with its $\mu$ determined for a given $N$ by
$N=\sum_{p}f_{p}^{-},$ (51)
which no longer stands for electroneutrality. Then, $U(N-1)-U(N)$ may be considered as a crude estimate of the thermal average of the first ionization energies. Likewise, $U(N)-U(N+1)$ is an approximate thermal first electron-attachment energy. At $T=0$, they correspond to the energy (the sign reversed) of the highest-occupied and lowest-unoccupied molecular or DysonOrtiz (2020) orbitals (HOMO and LUMO), respectively, according to Koopmans’ theoremSzabo and Ostlund (1982) or MBGF.Hirata et al. (2017) Since $U(N)$ is defined and computed with $\epsilon_{p}$, these internal-energy differences may furnish a physical meaning of the latter, which is indirect, yet potentially useful in relation to observables.
We call $U(N)-U(N-1)$ the thermal ionization energy and $U(N+1)-U(N)$ the thermal electron-attachment energy. Note the uncommon sign convention.
@@
The $U(N)$ thus defined are computed for an ideal gas of identical hydrogen fluoride molecules by thermal FCI, thermal QP(2), and thermal HF theories in Figs. 7, 8, and 9, respectively. The thermal ionization energies $U(N)-U(N-1)$ of these three methods are compared with one another and also with the thermal HOMO energies in Table 5. The same data for electron-attachment energies and LUMO are compiled in Table 6.
@@
Figure 7: Internal energy of thermal FCI theory $U_{\text{FCI}}$ (in $E_{\text{h}}$) of the hydrogen fluoride molecule (0.9168 Å) in the STO-3G basis set as a function of the average number of electrons $N$ relative to the value for $N=\bar{N}=10$.
Figure 8: Same as Fig. 7 but for the internal energy of thermal QP(2) theory $U_{\text{QP(2)}}$.
Figure 9: Same as Fig. 7 but for the internal energy of thermal HF theory $U_{\text{HF}}$.
@@
Table 5: Thermal ionization energy (in $E_{\text{h}}$) of the hydrogen fluoride molecule (0.9168 Å) in the STO-3G basis set as a function of temperature ($T$).
Thermal HOMO energy $U(N)-U(N-1)$
@@
$T$ / K QP(1)=HF [^1] First-order thermal quasi-particle theory or thermal HF theory. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference. QP(1)=HF [^1] First-order thermal quasi-particle theory or thermal HF theory. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference. FCI [^3] Thermal FCI theory.Kou and Hirata (2014)
@@
$10^{5}$ $-0.45147$ $-0.46765$ $-0.21004$ $-0.18425$ $-0.32041$
$10^{6}$ $-0.57384$ $-0.58577$ $-0.57181$ $-0.57425$ $-0.77028$
$10^{7}$ $-0.69361$ $-0.69692$ $-3.40146$ $-3.40165$ $-3.65153$
$10^{8}$ $-0.76988$ $-0.77209$ $-4.95141$ $-4.95143$ $-5.34456$
Table 6: Thermal electron-attachment energy (in $E_{\text{h}}$) of the hydrogen fluoride molecule (0.9168 Å) in the STO-3G basis set as a function of temperature ($T$).
Thermal LUMO energy $U(N+1)-U(N)$
$T$ / K QP(1)=HF [^1] First-order thermal quasi-particle theory or thermal HF theory. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference. QP(1)=HF [^1] First-order thermal quasi-particle theory or thermal HF theory. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference. FCI [^3] Thermal FCI theory.Kou and Hirata (2014)
@@
$10^{5}$ $0.48080$ $0.49392$ $0.07823$ $0.08932$ $0.40988$
$10^{6}$ $0.28118$ $0.20431$ $-0.55009$ $-0.55537$ $-0.12383$
$10^{7}$ $0.23384$ $0.13982$ $-3.14063$ $-3.14106$ $-2.71365$
$10^{8}$ $0.21118$ $0.11357$ $-4.90424$ $-4.90428$ $-4.48208$
From the tables, we observe that while the HOMO and LUMO energies of thermal quasi-particle theories correctly reduce to the ionization and electron-attachment energies of MBGF in the $T=0$ limit, they clearly have nothing to do with the approximate ionization or electron-attachment energies (as defined by $U(N)-U(N-1)$ or $U(N+1)-U(N)$ of thermal FCI theory) at $T\gg 0$, notwithstanding the questionable validity of the approximation. However, $U(N)-U(N-1)$ or $U(N+1)-U(N)$ calculated with $\epsilon_{p}$ of thermal quasi-particle theories are in line with the approximate ionization or electron-attachment energies of thermal FCI theory.
Thermal QP(2) theory is much closer to thermal FCI theory than thermal HF theory at low temperatures ($T\leq 10^{4}$ K), correctly accounting for electron correlation in thermal ionization and electron-attachment energies. However, at higher $T$, thermal HF and QP(2) theories are essentially the same, suggesting that the effect of electron correlation is surpassed by the lack thereof in the entropy and chemical potential in the quasi-independent-particle picture.
The figures reinforce these observations, but provide the following additional insights: As $T$ is lowered, the $U(N)$ curve begins to display the derivative discontinuity at $N=\bar{N}$, although the strict discontinuity does not occur until $T=0$. In both the low- and high-$T$ extremes, $U(N)$ has near-linear dependence on $N$ within each interval between adjacent integers, but at intermediate $T$, $U(N)$ is a convex curve with its minimum occurring away from $N=\bar{N}$.
Let us consider the slope of $U(N)$ at $N=\bar{N}$, which is another macroscopic quantity evaluable with $\epsilon_{p}$.
$\frac{\partial U}{\partial N}$ $=$ $\frac{\partial\mu}{\partial N}\sum_{p}\frac{\partial f_{p}^{-}}{\partial\mu}\frac{\partial U}{\partial f_{p}^{-}}$ (52)
$=$ $\frac{\sum_{p}f_{p}^{-}f_{p}^{+}\epsilon_{p}}{\sum_{p}f_{p}^{-}f_{p}^{+}},$
@@
$\frac{\partial f_{p}^{-}}{\partial\mu}=\beta f_{p}^{-}f_{p}^{+}.$ (53)
This is the increase in the internal energy upon infusion of an infinitesimal fraction of an electron into every molecule in the ideal gas. It is a weighted average of $\epsilon_{p}$ over all spinorbitals with the $p$th spinorbital weight being $f_{p}^{-}f_{p}^{+}$. It corresponds neither to the ionization nor to the electron-attachment process because it is oblivious to the sign of the variation in $N$. However, in the $T=0$ limit, it distinguishes these two processes by exhibiting the derivative discontinuity, i.e.,
$\lim_{T\to 0}\left(\frac{\partial U}{\partial N}\right)_{N=\bar{N}^{-}}$ $=$ $\epsilon_{\text{HOMO}},$ (54)
$\lim_{T\to 0}\left(\frac{\partial U}{\partial N}\right)_{N=\bar{N}^{+}}$ $=$ $\epsilon_{\text{LUMO}}.$ (55)
Until the $T=0$ limit is strictly reached, however, the $U(N)$ curve is smooth everywhere and the slope approaches the midpoint of the HOMO and LUMO energies. For $T\approx 0$,
$\left(\frac{\partial U}{\partial N}\right)_{N=\bar{N}}$$\approx$ $\frac{\epsilon_{\text{HOMO}}+\epsilon_{\text{LUMO}}}{2},$ (56)
which is also equal to the $T=0$ limit of the $\mu$.Hirata and Jha (2019); Hirata (2021b)
Table 7: $\partial U/\partial N$ ((in $E_{\text{h}}$)) of of the hydrogen fluoride molecule (0.9168 Å) in the STO-3G basis set as a function of temperature ($T$).
$T$ / K QP(1)=HF [^1] First-order thermal quasi-particle theory or thermal HF theory [Eq. (52)]. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference [Eq. (52)]. FCI [^3] Thermal FCI theory (numerical differentiation).Kou and Hirata (2014)
00444The average of ionization and electron-attachment energies of the corresponding MBGF.Hirata et al. (2017, 2024) $0.08253$ $0.12433$ $0.12371$

### v2

@@ III.1 Thermodynamic functions
$10^{5}$ $-101.02137$ $-103.48646$ $-101.30202$ $-102.10659$
$10^{6}$ $-150.56294$ $-151.43748$ $-150.60284$ $-151.24440$
$10^{7}$ $-729.93806$ $-730.10421$ $-729.94666$ $-730.09519$
$10^{8}$ $-6846.98049$ $-6847.00261$ $-6846.98165$ $-6847.00247$
Figure 3: The deviation from thermal FCI theory in grand potential $\Omega$ of an ideal gas of the identical hydrogen fluoride molecules (0.9168 Å) in the STO-3G basis set as a function of temperature ($T$). The “HF” stands for thermal HF theory, “MBPT(2)” denotes second-order finite-temperature many-body perturbation theory with the zero-temperature HF reference, “QP(2)” designates second-order thermal quasi-particle theory with the zero-temperature HF reference.
@@
00555Ground-state energy from zero-temperature HF, MBPT(2), and FCI theories. $-98.57076$ $-98.58809$ $-98.58809$ $-98.59658$
$10^{4}$ $-98.57076$ $-98.58809$ $-98.58809$ $-98.59658$
$10^{5}$ $-97.94385$ $-97.86604$ $-97.97596$ $-98.04938$
$10^{6}$ $-96.79410$ $-96.99284$ $-96.80270$ $-96.94534$
$10^{7}$ $-92.02773$ $-92.05724$ $-92.02910$ $-92.05557$
$10^{8}$ $-88.48266$ $-88.48744$ $-88.48288$ $-88.48740$
@@
Tables 14 list the thermodynamic functions—grand potential $\Omega$, internal energy $U$, chemical potential $\mu$, and entropy $S$—of thermal HF [thermal QP(1)],Gu and Hirata (2024) thermal QP(2), fintie-temperature MBPT(2),Hirata and Jha (2020); Hirata (2021a) and thermal full-configuration-interaction (FCI) theoriesKou and Hirata (2014) for an ideal gas of the identical hydrogen fluoride molecules in a wide range of temperature ($T$). Figures 36 plot the deviations from thermal FCI benchmarks, the latter being exact within a basis set. The thermal QP(2) and finite-temperature MBPT(2) calculations were based on the zero-temperature HF reference. See Ref. Gu and Hirata, 2024 for the data for the same system from Fermi–Dirac [thermal QP(0)] theory and other thermal mean-field theories, which do not take into account electron correlation.
In all cases, the thermal QP(2) results are close to finite-temperature MBPT(2) ones at low $T$ and are more accurate than thermal HF theory. This is expected because both thermal QP(2) theory and finite-temperature MBPT(2) reduce to zero-temperature MBPT(2) for internal energy, accounting for electron correlation, while thermal and zero-temperature HF theories do not. At higher $T$, thermal QP(2) theory quickly approaches thermal HF theory, where their quasi-independent-particle picture becomes relatively less erroneous. However, at $T=10^{6}$ K and $10^{7}$ K, finite-temperature MBPT(2) is systematically more accurate as it accounts for the effect of electron correlation in $S$ and $\mu$ and it uses more complete second-order corrections (including anomalous-diagram terms) in $\Omega$ and $U$. At the intermediate temperature of $T=10^{5}$ K, thermal QP(2) theory (or even thermal HF theory) outperforms finite-temperature MBPT(2). This may be ascribed to the fact that thermal QP(2) theory includes the effect of electron correlation in the orbital energies, and thus higher-order correlation corrections in a manner similar in spirit to self-consistent Green’s function methods,Luttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Van Neck, Waroquier, and Ryckebusch (1991); Dickhoff (1999); Dickhoff and Barbieri (2004); Dahlen and van Leeuwen (2005); Barbieri (2006); Barbieri and Dickhoff (2009); Phillips and Zgid (2014); Neuhauser, Baer, and Zgid (2017); Tarantino et al. (2017); Kadanoff and Baym (2018); Coveney and Tew (2023) which are said to work for strong correlation. However, this similarity must not be overstated because self-consistent Green’s function methods replace reference orbitals and orbital energies by electron-correlated counterparts, while thermal QP(2) theory uses correlated orbital energies only in the Fermi–Dirac distribution functions and thus reduces identically to MBPT(2) at $T=0$.
Additional numerical data are presented in Supplementary Information,sup which reinforces the foregoing conclusions.
III.2 Thermal quasi-particle energies
@@
$\epsilon_{p}^{\text{QP}}\equiv\frac{\partial U^{\text{QP}}}{\partial f_{p}^{-}},$ (46)
which encompasses all of the Fermi–Dirac, thermal HF, and thermal QP(2) ansätze.
Hence, the $\epsilon_{p}^{\text{QP}}$ has the literal physical meaning of the increase in the internal energy upon infusion of an infinitesimal fraction of an electron in the $p$th spinorbital. This interpretation can be viewed as a finite-temperature version of Janak’s theoremJanak (1978) in DFT, and may be called thermal Janak’s theorem. It is not unreasonable to consider a fraction of an electron here because the process in question is the thermal average of the same processes involving an infinite number of particles.
While the $\epsilon_{p}^{\text{QP}}$ at $T=0$ is rigorously related to the ionization and electron-attachment energies of a molecule or solid according to Koopmans’ theoremSzabo and Ostlund (1982) or MBGF,Hirata et al. (2017) it cannot be expected to be even close to their thermal averages at $T\gg 0$ because the latter involve different orbitals at different probabilities, even in a strict independent-particle picture.
Therefore, the simplest reasonable approximations to thermal ionization ($I^{\text{QP}}$) and electron-attachment ($A^{\text{QP}}$) energies (signs reversed) based on the $\epsilon_{p}^{\text{QP}}$ may be their weighted averages such as
$I^{\text{QP}}$ $=$ $\sum_{p}\epsilon_{p}^{\text{QP}}f_{p}^{-}(\bar{N})-\sum_{p}\epsilon_{p}^{\text{QP}}f_{p}^{-}(\bar{N}-1),$ (47)
$A^{\text{QP}}$ $=$ $\sum_{p}\epsilon_{p}^{\text{QP}}f_{p}^{-}(\bar{N}+1)-\sum_{p}\epsilon_{p}^{\text{QP}}f_{p}^{-}(\bar{N}),$ (48)
where $\bar{N}$ is the average number of electrons that ensures electroneutrality and $f_{p}^{-}(N)$ is the Fermi–Dirac distribution function for an arbitrary (integer or noninteger) average number of electrons $N$, i.e.,
$N=\sum_{p}f_{p}^{-}(N).$ (49)
It should be understood that only the chemical potential $\mu^{\text{QP}}$ (but not $\epsilon_{p}^{\text{QP}}$) is varied so that this equation is satisfied.
We examine the validity of thermal Janak’s theorem by comparing the above approximations ($I^{\text{QP}}$ and $A^{\text{QP}}$) against $\Delta U^{\text{FCI}}$ defined by
$\Delta U^{\text{FCI}}$$\equiv$ $\left\{\begin{array}[]{ll}U^{\text{FCI}}(\bar{N})-U^{\text{FCI}}(\bar{N}-1),&\text{ionization};\\ U^{\text{FCI}}(\bar{N}+1)-U^{\text{FCI}}(\bar{N}),&\text{electron attachment},\end{array}\right.$ (52)
where $U^{\text{FCI}}(N)$ is the internal energy of thermal FCI theory for an ideal gas of identical molecules, each of which has $N$ electrons on average. It is
$U^{\text{FCI}}(N)$$\equiv$ $\frac{\sum_{I}E_{I}e^{-\beta(E_{I}-\mu N_{I})}}{\sum_{I}e^{-\beta(E_{I}-\mu N_{I})}},$ (54)
where $E_{I}$ is the FCI energy of the $I$th state and the chemical potential $\mu$ is determined by
$N$ $=$ $\frac{\sum_{I}N_{I}e^{-\beta(E_{I}-\mu N_{I})}}{\sum_{I}e^{-\beta(E_{I}-\mu N_{I})}}.$ (55)
@@
This is a valid and in fact, exact (within a basis set) treatment of thermodynamics if the particles are not electrically charged. For electrons, however, it describes a physically unrealistic, massively charged plasma, whose energy is not even extensive.Fisher and Ruelle (1966); Dyson and Lenard (1967); Hirata and Ohnishi (2012) Nevertheless, $U^{\text{FCI}}(N)$ and $\Delta U^{\text{FCI}}$ are computationally well defined even for electrons thanks to the ideal-gas assumption (i.e., no interactions between molecules) and will be useful for the following analysis. It may or may not be a reasonable approximation to the thermal average of ionization or electron-attachment energies.
The approximations of Eqs. (47) and (48) imply the corresponding internal-energy functional of the form,
$U^{\text{QP}}(N)\equiv\sum_{p}\epsilon_{p}^{\text{QP}}f_{p}^{-}(N),$ (56)
which leads to
$\Delta U^{\text{QP}}$$\equiv$ $\left\{\begin{array}[]{ll}I^{\text{QP}},&\text{ionization};\\ A^{\text{QP}},&\text{electron attachment}.\end{array}\right.$ (59)
@@
The $U^{\text{FCI}}(N)$ and $U^{\text{QP}}(N)$ thus defined are computed for an ideal gas of the identical hydrogen fluoride molecules by thermal FCI, thermal QP(2), and thermal HF theories and are plotted as a function of $N$ in Figs. 7, 8, and 9, respectively. The thermal ionization energies of these three methods are compared with one another and also with the thermal HOMO energies in Table 5. The same data for electron-attachment energies and LUMO are compiled in Table 6.
@@
Figure 7: Internal energy of thermal FCI theory $U^{\text{FCI}}$ (in $E_{\text{h}}$) of an ideal gas of the identical hydrogen fluoride molecules (0.9168 Å) in the STO-3G basis set as a function of the average number of electrons $N$ relative to the value at $N=\bar{N}=10$.
Figure 8: Same as Fig. 7 but for the internal energy of thermal QP(2) theory $U^{\text{QP(2)}}$.
Figure 9: Same as Fig. 7 but for the internal energy of thermal HF theory $U^{\text{HF}}$.
@@
Table 5: Thermal ionization energy (in $E_{\text{h}}$) of an ideal gas of the identical hydrogen fluoride molecules (0.9168 Å) in the STO-3G basis set as a function of temperature ($T$).
$\epsilon_{p}^{\text{QP}}$ $\Delta U$
@@
$T$ / K HF [^1] Thermal HF theory. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference. HF [^1] Thermal HF theory. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference. FCI [^3] Thermal FCI theory.Kou and Hirata (2014)
@@
$10^{5}$ $-0.45147$ $-0.41998$ $-0.21004$ $-0.18483$ $-0.32041$
$10^{6}$ $-0.57384$ $-0.57392$ $-0.57181$ $-0.56735$ $-0.77028$
$10^{7}$ $-0.69361$ $-0.69551$ $-3.40146$ $-3.40055$ $-3.65153$
$10^{8}$ $-0.76988$ $-0.77193$ $-4.95141$ $-4.95127$ $-5.34456$
Table 6: Thermal electron-attachment energy (in $E_{\text{h}}$) of an ideal gas of the identical hydrogen fluoride molecules (0.9168 Å) in the STO-3G basis set as a function of temperature ($T$).
$\epsilon_{p}^{\text{QP}}$ $\Delta U$
$T$ / K HF [^1] Thermal HF theory. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference. HF [^1] Thermal HF theory. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference. FCI [^3] Thermal FCI theory.Kou and Hirata (2014)
@@
$10^{5}$ $0.48080$ $0.50816$ $0.07823$ $0.10621$ $0.40988$
$10^{6}$ $0.28118$ $0.31458$ $-0.55009$ $-0.54444$ $-0.12383$
$10^{7}$ $0.23384$ $0.27782$ $-3.14063$ $-3.13959$ $-2.71365$
$10^{8}$ $0.21118$ $0.26168$ $-4.90424$ $-4.90408$ $-4.48208$
From the tables, we observe that while the HOMO and LUMO energies of thermal quasi-particle theories correctly reduce to the ionization and electron-attachment energies of MBGF in the $T=0$ limit, they clearly have nothing to do with the approximate thermal ionization or electron-attachment energies as defined by $\Delta U^{\text{FCI}}$ at $T\gg 0$, notwithstanding the questionable validity of the latter. However, $\Delta U^{\text{QP}}$ calculated with $\epsilon_{p}^{\text{QP}}$ are in line with the approximate thermal ionization or electron-attachment energies, giving credence to the physical meaning of thermal orbital energies provided by thermal Janak’s theorem.
Thermal QP(2) theory is much closer to thermal FCI theory than thermal HF theory at low temperatures ($T\leq 10^{4}$ K), correctly accounting for electron correlation in thermal ionization and electron-attachment energies. However, at higher $T$, thermal HF and QP(2) theories are essentially the same, suggesting that the effect of electron correlation is surpassed by the lack thereof in the entropy and chemical-potential terms that dominate at such temperatures.
The figures reinforce these observations, but provide the following additional insights: As $T$ is lowered, all $U(N)$ curves begin to display signs of a derivative discontinuity at $N=\bar{N}$, although the strict discontinuity does not occur until $T=0$. Recall that all nonhybrid DFT approximations lack a derivative discontinuity,Perdew and Levy (1983) causing a gross underestimation of band gaps. In both the low- and high-$T$ extremes, $U(N)$ has near-linear dependence on $N$ within each interval between adjacent integers, but at intermediate $T$, each $U(N)$ curve is convex with its minimum occurring away from $N=\bar{N}$.
Let us consider the slope of $U^{\text{QP}}(N)$ at $N=\bar{N}$, which is another macroscopic quantity evaluable with $\epsilon_{p}^{\text{QP}}$.
$\frac{\partial U^{\text{QP}}}{\partial N}$ $=$ $\frac{\partial\mu^{\text{QP}}}{\partial N}\sum_{p}\frac{\partial f_{p}^{-}}{\partial\mu^{\text{QP}}}\frac{\partial U^{\text{QP}}}{\partial f_{p}^{-}}$ (60)
$=$ $\frac{\sum_{p}f_{p}^{-}f_{p}^{+}\epsilon_{p}^{\text{QP}}}{\sum_{p}f_{p}^{-}f_{p}^{+}},$
@@
$\frac{\partial f_{p}^{-}}{\partial\mu^{\text{QP}}}=\beta f_{p}^{-}f_{p}^{+}.$ (61)
This is the increase in the internal energy upon infusion of an infinitesimal fraction of an electron into every molecule in the ideal gas. It is a weighted average of $\epsilon_{p}^{\text{QP}}$ over all spinorbitals. It corresponds to neither the ionization nor electron-attachment process because it is oblivious to the sign of the variation in $N$. However, in the $T=0$ limit, it distinguishes these two processes by exhibiting the derivative discontinuity, i.e.,
$\lim_{T\to 0}\left(\frac{\partial U^{\text{QP}}}{\partial N}\right)_{N=\bar{N}^{-}}$ $=$ $\epsilon_{\text{HOMO}}^{\text{QP}},$ (62)
$\lim_{T\to 0}\left(\frac{\partial U^{\text{QP}}}{\partial N}\right)_{N=\bar{N}^{+}}$ $=$ $\epsilon_{\text{LUMO}}^{\text{QP}}.$ (63)
Until the $T=0$ limit is strictly reached, however, the $U^{\text{QP}}(N)$ curve is smooth everywhere and the slope approaches the midpoint of the HOMO and LUMO energies.Perdew and Levy (1983) For $T\approx 0$,
$\left(\frac{\partial U^{\text{QP}}}{\partial N}\right)_{N=\bar{N}}$$\approx$ $\frac{\epsilon_{\text{HOMO}}^{\text{QP}}+\epsilon_{\text{LUMO}}^{\text{QP}}}{2},$ (64)
which is also equal to the $T=0$ limit of the $\mu^{\text{QP}}$.Hirata and Jha (2019); Hirata (2021b)
Table 7: $\partial U/\partial N$ (in $E_{\text{h}}$) of an ideal gas of the identical hydrogen fluoride molecules (0.9168 Å) in the STO-3G basis set as a function of temperature ($T$).
$T$ / K HF [^1] Thermal HF theory [Eq. (60)]. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference [Eq. (60)]. FCI [^3] Thermal FCI theory (numerical differentiation).Kou and Hirata (2014)
00444The average of the ionization and electron-attachment energies (signs reversed) of the corresponding MBGF.Hirata et al. (2017, 2024) $0.08253$ $0.12433$ $0.12371$